OpenStudy (anonymous):

Linda and Dave leave simultaneously from the same starting point biking in opposite directions. Linda bikes at 7 miles per hour and Dave bikes at 9 miles per hour. How long will it be until they are 24 miles apart from each other?

OpenStudy (anonymous):

7t+9t=24 t=24/15 h

OpenStudy (radar):

t=24/16 h

OpenStudy (anonymous):

right